OpenStudy (hedgehog):

Write an equation for the line that contains the point (-6, -8) and is parallel to the line 3x – 2y = -12.

OpenStudy (anonymous):

Since line is parallel to the line 3x – 2y = -12. The eq. 3x – 2y = -12 can be written as 2y = 3x + 12 i.e. \[i.e. y = \frac{ 3 }{ 2 } x + 6\] Now comparing it with y = mx + c we have slope \[m=\frac{ 3 }{ 2 }\] Since slopes of parallel lines are same, therefore slope required line will also be \[m=\frac{ 3 }{ 2 }\] Now since line passes through the points (-6, -8) i.e. x1 = -6, y1 = -8 Eq of line in slope point form is given as y-y1=m(x-x1) i.e. y-(-8) =3/2[x-(-6)] i.e. y+8 =3/2[x+6] i.e. 2 y+16 =3x+18 i.e. 3x -2y +18-16=0 i.e. 3x -2y +2=0 and this the required eq of line

OpenStudy (anonymous):

Thank you. Good Night friends